OpenStudy (anonymous):

What is the molarity of an NaOH solution if 3.47 mL is titrated by 11.1 mL of 0.0904 M HNO3?

OpenStudy (somy):

write a reaction between NaOH and HNO3

OpenStudy (somy):

make sure balance it

OpenStudy (abmon98):

NaOH(aq) + HNO3(aq) --> NaNO3(aq) +H2O(l) number of moles=concentration*volume 11.1ml=0.0111dm^3 c=0.0904mol/dm^3, 0.0111*0.0904= 0.00100344 moles of HNO# ratio of HNoratio of HNO3 to NaOH is 1:1 0.00100344=0.00347*concentration concentration=0.00100344/0.00347=0.289mol/dm^3